OpenStudy (anonymous):

solve x(squared) - y(squared)=7 y=1+x/2

OpenStudy (anonymous):

here's an online equation graph, type in 1/x^2 where it says y= then press draw and it'll graph the equation. If you have a graphing calculator you can also use that to graph it. http://www.freemathhelp.com/equation-graph

OpenStudy (amorfide):

x²-y²=7 y=1+x/2 you are given what y is substitute y into x²-y²=7

OpenStudy (amorfide):

x²-y²=7 x²-(1+x/2)²=7 can you do it from here?

OpenStudy (anonymous):

i hv done it but couldnt finish it

OpenStudy (anonymous):

cuz the answer i m gettin is wrong

OpenStudy (anonymous):

can u please solve it fully for me thanx

OpenStudy (amorfide):

okay